the crystalline salt Na2So4.xH2O on heating loses 55.9%of its mass and becomes anhydrous .the formula of crystlline salt

Grade:12th pass

1 Answers

Arun
4 years ago
Molar mass of salt = (2 × 23) + 32 + (16 × 4) + x × 18
= 142 + 18x
On heating it will lose water and become anhydrous. 55.9% mass is the mass of water in Na2SO4.xH2O.
Mass by mass% = (mass / molar mass of compound) × 100
mass of water = 18x
55.9 = (18x / 142 + 18x) × 100
55.9(142 + 18x) = 1800x
7937.8 + 1006.2x = 1800x
x = 9.99
x = 10 (approx.)
So, the molecular formula of compound is Na2SO4.10H2O.